OpenStudy (anonymous):

Suppose 5 men and 7 women are on a crowded elevator. At the next floor, four people get off the elevator. Find the probability that two are women.

OpenStudy (anonymous):

And find the probability that one is a woman.

OpenStudy (zubhanwc3):

to the one one, its 7/13

OpenStudy (anonymous):

Thats not an answer for either

OpenStudy (zubhanwc3):

argh, havnt done this in 4 years, and im racking my brain:P gimmie a moment,

OpenStudy (anonymous):

okay thanks, ive just been doing all sorts of stuff and i got nothing

OpenStudy (anonymous):

This is sampling without replacement. Seems complicated. Might find the number of combinations of 12 things taken four at a time. Perhaps looking at those foursomes would be productive. I don't now.

OpenStudy (anonymous):

if you meant exactly two women then (7C2) (5C2) / (12C4) = 14/33

OpenStudy (anonymous):

yea that was it i think

OpenStudy (anonymous):

do you mind helping me with the "one woman" part also? @sourwing

OpenStudy (anonymous):

one women means, 1W and 3M. (7C1) (5C3) / (12C4) = 14/99

OpenStudy (anonymous):

Thanks for explaining too i was not getting it, but i got it now! Thanx! @sourwing

OpenStudy (anonymous):

@sourwing Very nicely done.